Diffusion Morning Report June 30, 2004

Ray Seymour seymour at snf.stanford.edu
Wed Jun 30 10:08:12 PDT 2004


Morning all,
BPSG and Nitride are both down for pull and clean. tube 1 was ran
yesterday afternoon and no problems found, but I will leave the problem
on it for a while. Epi HCL to be changed in a little while. All else is
looking good at this time.


Diffusion                   Date: 6/30/2004
Tube Name/Number  Status
Comments                               THICKNESS-Change at
Tylan 1 problem                                                    Ran
ok yesterday afternoon
Tylan 2 problem
Tylan 3 Up
Tylan 4 Up
Tylan 5-BBr3 Up
Tylan 6-POCl3 Up
Tylan 7-TEOS2 down                                           Qual
testing-recipe issues
Tylan 8-SiGe
Up
8.57/30
Tylan 9-Poly
Up
10.82/30
Tylan 10-Nitride Down                                          Pull and
clean                              0/30
Tylan 11-BPSG-LTO Up                                       Pull and
clean                              0/30
Tylan 12-FGA Up
Tystar
Up
13.5/30
Doped Poly GONE
TEOS yellow  running
Tylan 16 N/A
Burn Box Up
ZYGO Up

GAAS 22 UP
GAAS 23 UP

SiH4 for furnaces                             Auto changeover working
well  A on line 360 PSI
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040630/5d8e9fd5/attachment.html>


More information about the diffusion mailing list